If tantita = 7 / 8 find sintita + costita by sintita-costita

Grade:10

1 Answers

Sharmila
6 years ago
We know that tan A = Sin A/cosA =7/8Applying componendo and dividendo if a/b=c/d then a+b/a-b =c+d/c-d,The ans will be 7+8/7-8=-15